An unknown force accelerates a 6500kg moving van from -3m/s to -13m/s in a time of 2s. In Newtons, how large is the force?


1
Expert's answer
2021-01-29T01:28:28-0500

By definition, the acceleration is:


"a = \dfrac{v_f-v_i}{t}"

where "v_f = -13m\/s" is the final velocity, "v_i = -3m\/s" is the initial velocity, "t=2s" is the time of acceleration.

According to the second Newton's law, the force acting of the van is then:


"F = ma = m\dfrac{v_f-v_i}{t}"

where "m = 6500kg" is the mass. Thus, obtain:


"F=6500\cdot\dfrac{-13-(-3)}{2} = -52000N"

Answer. 52000 N.